OpenStudy (anonymous):

compare the numbers in each exercise using an inequality symbol. 5 2/3, sqrt 29

OpenStudy (anonymous):

5 2/3 =17/3 = sqrt 32.111 sqrt 29 for f(x) = sqrt (x), it is decreasing function therefore, sqrt 32.111 > sqrt 29 5 2/3 > sqrt 29